Here is part of the table of a linear function. Figure out an equation for the
function.
x y
4 17
5 19
Explain in words how you determined the equation.

Answer by stanbon(75887)

Here is part of the table of a linear function. Figure out an equation for the
function.
x y
4 17
5 19
Explain in words how you determined the equation.
----------------------
Form: y = mx + b
slope = (19-17)/(5-4) = 2
----
Substitute for y,m,x and solve for "b":
17 = 2*4 + b
b = 9
------
Equation:
y = 2x + 9
================
Cheers,
Stan H.
================